### v1

Abstract
Real-world development demands code that is readable, extensible, and testable by organizing the implementation into modular components and iteratively reuse pre-implemented code. We term this iterative, multi-turn process codeflow and introduce CodeFlowBench, the first benchmark designed for comprehensively evaluating LLMs’ ability to perform codeflow, namely to implement new functionality by reusing existing functions over multiple turns. CodeFlowBench comprises 5,258 problems drawn from Codeforces and is continuously updated via an automated pipeline that decomposes each problem into a series of function-level subproblems based on its dependency tree and each subproblem is paired with unit tests. We further propose a novel evaluation framework with tasks and metrics tailored to multi-turn code reuse to assess model performance. In experiments across various LLMs under both multi-turn and single-turn patterns. We observe models’ poor performance on CodeFlowBench, with a substantial performance drop in the iterative codeflow scenario. For instance, o1-mini achieves a pass@1 of 20.8% in multi-turn pattern versus 37.8% in single-turn pattern. Further analysis shows that different models excel at different dependency depths, yet all struggle to correctly solve structurally complex problems, highlighting challenges for current LLMs to serve as code generation tools when performing codeflow. Overall, CodeFlowBench offers a comprehensive benchmark and new insights into LLM capabilities for multi-turn, iterative code generation, guiding future advances in code generation tasks. The code and dataset will be available soon on https://github.com/IAAR-Shanghai/CodeFlowBench
1 Introduction
Large Language Models (LLMs) have been widely applied to code generation tasks and many benchmarks have been developed to evaluate LLMs’ performance on code generation tasks, such as HumanEval[3], MBPP[2], and DS-1000[15]. With the rapid advancement of LLMs’ capabilities, the focus of using LLMs for code generation has shifted from merely tackling toy problems to integrating these models into real-world development workflows for more complex code implementations with the rapid advancements of the capabilities of LLMs.Correspondingly, several benchmarks have begun to ground their evaluation content in practical development scenarios. For instance, BigCodeBench[29] selects functions directly from open-source repositories and requires models to invoke external libraries while SWE-Bench[13] evaluates bug-fixing capabilities via GitHub pull requests.
However, we contend that current benchmarks still fail to capture two fundamental and pivotal aspects of real-world development workflows: multi-turn and iterative. To enhance the readability, extensibility and testability of code, developers tend to organize their workflow as a codeflow where they start by implementing the lowest-level functions to solve fundamental subproblems and then reuse these functions to build more sophisticated modules. An example of codeflow is shown in Figure 1. Codeflow is inherently a multi-turn, iterative code generation process since developers refine or extend functionality by leveraging pre-implemented code in each turn.
Figure 1: An example of building a currency arbitrage detection and execution system. The implementation in left side is organized in a codeflow format where the developer first creates three modular functions and then reuses them in the main function to process input data and determine optimal strategies. This approach allows for easier testing, debugging and extension of each individual component. In contrast, the right side shows a single-turn implementation where the entire process is combined into one block, making it harder to read, test and maintain.
To serve as an excellent code-generation tools in a codeflow, LLMs must handle developers’ multi-turn requests, understand the dependency relationships between existing code and newly requested code in each turn and then produce reusable code for subsequent iterations. However, we find that existing benchmarks like MTPB[21] and CodeEvalPro[28] lack sufficient difficulty and do not provide a comprehensive, fine-grained evaluation framework for LLMs’ multi-turn, iterative code generation capabilities. Therefore, we believe there remains a need for a more challenging benchmark that can thoroughly evaluate LLMs’ capabilities in a codeflow.
To fill the gap, we introduce a new evaluation benchmark called CodeFlowBench. Unlike benchmarks mentioned above, CodeFlowBench sources its problems from the competitive programming platform Codeforces [^1] https://codeforces.com/, which offers several advantages:(1) High difficulty. Even state-of-the-art LLMs have low pass rates on Codeforces problems[22]. Compared to simple and straightforward code implementation, competitive problems can make multi-turn, iterative solution process more realistic and meaningful. (2) High quality. Unlike many open-source repositories, Codeforces publishes official editorial and provides automated correctness checks, ensuring the integrity of our dataset’s problems.(3) Real-time updates. To maintain uncontaminated test data and ensure fairness, it is essential that benchmarks can be refreshed regularly. Codeforces releases new problems at a high frequency, enabling continuous updates for CodeFlowBench.We also developed an automated, lightweight pipeline that processing problems in Codeforces to be sets of multi-turn, iterative subproblems to build CodeFlowBench and make it available for continuous updates.
In this paper, we introduce the initial version of CodeFlowBench, which comprises 5,258 test problems extracted from the Codeforces problem archive accumulated until now. To ensure ongoing fairness and freshness, we will periodically run our automated pipeline and continuously update the benchmark with new, uncontaminated problems.
In sum, the contribution of this paper is threefold:
  • We build a data curation pipeline based on the Codeforces platform. This pipeline decomposes existing solutions based on function dependencies to form multi-turn, iterative code problems that require code reuse. The approach is fully automated, lightweight, and easy to maintain, allowing for real-time updates.
  • We introduce the first benchmark to specifically target the multi-turn, iterative code generation capabilities of LLMs. It features a large and diverse collection of competition-level code problems. By leveraging official editorial solutions, automatically verifying generated code and capturing test cases through input-output streams, we ensure the robustness and quality of our dataset.
  • We propose a novel evaluation task framework that incorporates both multi-turn and single-turn code generation patterns. We also introduce new structural labels and metrics derived from the dependency tree of each problem, aiming to more comprehensively capture the characteristics of multi-turn tasks and bridge the gap between existing evaluation metrics and the unique challenges posed by multi-turn tasks. The experimental findings reveal important characteristics of model capabilities and deficiencies in multi-turn, iterative code generation, offering meaningful directions for future research and development.
2 Related Work
Code Generation Benchmarks
The landscape of code generation benchmarks has been progressing from simpler to more complex tasks. Pioneering works like HumanEval[3] and MBPP[2] center on code generation for stand-alone function. The main constraints are simple problem difficulty and limited dependency environments. Building on these works, recent benchmarks have emerged to evaluate various aspects of LLMs as code generation tools: Benchmarks such as APPS[7], CodeContest[17], LiveCodeBench[11], and CodeElo[22] employ complex competition-style problems to rigorously test the models’ ability to understand and reason about code; Benchmarks like MTPB[21], InterCode[27] and TreeInstruct[14] focus on the iterative, multi-turn dialogue approach users often adopt when generating code with LLMs, thereby evaluating the models’ ability to retain context, understand, and reason across multiple interactions. Meanwhile, CodeEvalPro[28] uses existing functions to create self-invoking functions, testing the models’ ability to solve these tasks by invoking existing code. Besides, LiveCodeBench[11] and EvoCodeBench[16] plan regular updates to their problem sets, aiming to mitigate the risk of data leakage and maintain testing fairness. Collectively, these benchmarks significantly contribute to the development and refinement of evaluation frameworks. Our work builds upon these cutting-edge benchmarks, focusing on the multi-turn, iterative code generation needs of developers tackling complex implementations, and offering a novel, high-quality benchmark to address these challenges.
Code Generation LLMs
In recent years, large language models have achieved remarkable and rapid progress in code generation. Early works such as Codex[3] and AlphaCode[17] have demonstrated impressive performance in a wide range of tasks from simple function completion to solving competition-level programming problems. With the scaling up of pre-trained models, exemplified by GPT-4[1], Code-Llama[23], Deepseek-Coder[5], and Qwen2.5-Coder[9], advanced pre-trained models now have high-level performance across various programming tasks in different languages and domains. Building on these foundations, the code generation capabilities of LLMs have further advanced through instruction tuning and agent frameworks. Instruction tuning Models like WizardCoder[19] and MagicCoder[25] exhibit stronger interactive dialogue capabilities and a better grasp of user intent, while agent-based frameworks such as Agentcoder[8] and Mapcoder[10] enhance the autonomy of LLMs in code generation and enable iterative planning, execution, and self-evaluation in code generation task. Overall, LLMs are rapidly becoming indispensable tools for developers, significantly reducing the time and effort required to produce high-quality code.[12]. Our work provide a new evaluation approach and direction for the advancement of LLMs’ code generation capabilities by targeting the multi-turn, iterative feature of code generation.
3 CodeFlowBench
CodeFlowBench is a lively-updated benchmark which contains 5,258 diverse, high-quality and challenging problems.An example problem of CodeFlowBench is shown in Figure 2. Remaining part will introduce the data curation pipeline and the evaluation method of CodeFlowBench.
Figure 2: An example problem from CodeFlowBench, which is originally sourced from problem 2055E in Codeforces.The top-left block displays the whole problem information, including scraped data as well as the designated overall-depth and overall-turns metrics. The bottom-left block illustrates the dependency structure of the problem, where each node represents a subproblem and each connecting line indicates a dependency relationship. The right block details each subproblem. Models are required to engage in this codeflow by solving the subproblems iteratively, turn by turn. The bottom block shows the LLM interaction workflow for each turn.
3.1 Data Curation Pipeline
We developed an automated, lightweight data-curation pipeline for generating multi-turn coding problems. The pipeline consists of two main stages: (1)Scraping and processing data (Stages I–III): it contains source-specific routines that collect and normalize raw input from various repositories. (2)Decomposing and generating subproblems (Stages IV–VI): it contain a general framework that, treats each function as an individual subproblem and decomposes the problem according to the functions’ dependency relationships given a codebase.
It is the first pipeline to incorporate the official editorials of Codeforces problems and to include a submission bot for fair verification across all Codeforces-targeted pipelines. By incorporating the official editorials, we avoid scraping user solutions which may raise ethical concern, eliminating the need for extensive filtering and ensure that the codeflow for each problem is organized within an optimal framework, thereby preserving the high quality of our benchmark. We believe our pipeline can also contribute to future competitive code generation benchmarks. Specifically, our pipeline comprises the following steps.
Figure 3: Illustration of the data curation pipeline of CodeFlowBench
Stage I: Scrape Problem Information
We visit the Codeforces problem list [^2] https://codeforces.com/problemset and then navigate to each problem’s page. We extract the following information: problem ID, title, time and memory complexity limits, problem description, input description, output description, sample tests, notes. Additionally, by querying the official Codeforces API, we retrieve each problem’s rating and tags. The detailed explanation of the information and an example can be found in Appendix A.1.
Stage II: Scrape Editorial Information
We visit the tutorial page from problem website. Since tutorials are typically collections of problems grouped by a contest round, we opted to scrape explanations at the round level. The explanation pages display each problem’s identifier, textual explanation and code solution in sequence. Therefore, we first locate each problem’s identifier and extracts the content between adjacent identifiers as the solution content of corresponding problem. We primarily rely on the URL corresponding to each problem ID in the HTML code for identifier recognition, supplemented by common textual patterns to enhance accuracy. Once the scraping is complete, we map these identifiers to the problem IDs collected in Stage I to append the explanations to each problem. Additionally, we filter out explanations of evidently low quality and split the explanations into text and code segments to facilitate model processing. A detailed explanation of the method and an example can be found in Appendix A.2
Stage III: Solution Code Generation
We feed both the problem and its explanation information into the Deepseek-R1 model[6], allowing it to either directly use the code implementation if provided or organize the existing textual solution and convert it to parsable solution code that meets the requirements for extracting function dependencies and constructing subproblems. The generated code is then submitted to the Codeforces platform for correctness verification, and we retain only those codes that successfully solve the problem. The prompt template can be found in Appendix A.3
Stage IV: Subproblems Generation
We first parsing code generated in Stage III with its Abstract Syntax Tree(AST) to obtain its function dependency relationship. To select potential subproblem and correctly arrange it as a codeflow, we filter out low-quality functions and applying topological sort based on dependencies to ensure that lower-level functions precede higher-level ones. Each filtered function is treated as a subproblem. For every function, we extract its corresponding code segment from the solution and record its depth in the AST along with the functions it depends on. We then input this information, together with the problem and explanation details, into the Deepseek-V3 model[18] and require the model generates a problem description for each function implementation, thereby forming a complete subproblem that contain name, statement, depth, dependencies, solutions attributes.The prompt template and an example can be found in Appendix A.4
Stage V:Test Cases Generation
For each problem, we first scrape the publicly available test cases from the Codeforces platform and use these as inputs to execute the corresponding solution code. Meanwhile, we embed input-output capture logic at every location where a function is invoked within the solution code. During execution, we record the input-output pair for each function call. These input-output pairs serve as test cases for each subproblem. We explain the process in detail in Appendix A.5.
Stage VI: Automatic Labeling
For each problem, we define two metrics as labels to indicate the complexity of the problem’s structure, which are useful for our experiment analysis.
  • Overall-Turns: The number of turns required to solve the entire problem.
  • Overall-Depth: The depth of the AST. Each subproblem serves as a node in the tree, with the main/solve function as the root node.
The distribution of overall-turns and overall-depths in CodeFlowBench is shown in 4(a). We also compare the Rating metric provided by Codeforces platform which reflects the difficulty of a problem with our metrics.The strong correlation shown in 4(b) confirms that our label effectively captures the intrinsic complexity of these problems.
Figure 4: Statistics of the overall-turns and overall-depth metrics in CodeFlowBench. Subfigure (a) illustrates the distribution of these two metrics. Subfigure (b) depicts their correlation with the Rating metric.
3.2 Evaluation Framework
Partially like the task defined in HumanEval[3], models evaluated by CodeFlowBench are also required to complete the code implementation for a given function in each turn. The main difference lies in the supporting materials of the task. In addition to the function signature and problem description, we also provide dependency information and pre-implemented functions. We require models to reuse existing code if needed. Moreover, we provide an overall background description of the whole problem since we consider the opportunistic scenario where developers have knowledge of the whole problem as a ultimate goal when implementing code in a turn to be more common and reflective of real-world situations.
The evaluation task can be formally defined. Suppose $F_{t}$ is the function for which the implementation $C_{t}$ is to be generated at turn $t$ and $S_{t}$ is the corresponding problem statement for $F_{t}$, which may include the description of inputs, outputs, and functionality of $F_{t}$. $\mathcal{D}=\{k_{1},k_{2},\ldots,k_{d}\}$ denote the set of previous turns on which the current turn depends and $F_{k_{1}},F_{k_{2}},\ldots,F_{k_{d}}$ represent the definitions of dependency function. $B$ is the background information for the overall problem and $C_{1},C_{2},\ldots,C_{t-1}$ denote the implementations generated in the chatting history. Suppose the total number of turns is $T$. Then the implementation $C_{t}$ at turn $t$ is generated as follows $(t=1,2,\cdots,T)$:
$C_{t}\leftarrow p\Bigl{(}\cdot\,\Bigm{|}\,F_{t},S_{t},\{F_{k_{1}},F_{k_{2}},\ldots,F_{k_{d}}\},\{C_{1},C_{2},\ldots,C_{t-1}\},B\Bigr{)}.$
For comparison, we also provide a single-turn evaluation pattern to assess the model’s multi-turn capabilities. In the single-turn task, the final implementation is generated by:
$\{C_{1},C_{2},\cdots,C_{T}\}\leftarrow p\Bigl{(}\cdot\,\Bigm{|}\,\{F_{1},F_{2},\cdots,F_{T}\},\{S_{1},S_{2},\cdots,S_{T}\},B\Bigr{)}.$
Pass@k [3] is a widely used evaluation metric for code generation tasks. We adopt it as our primary evaluation metric for both multi-turn and single-turn evaluation patterns. However, a notable limitation of using $pass@k$ in a multi-turn setting is that it only provides a final, coarse-grained evaluation and cannot fully distinguish the capabilities of different models. In other words, pass@k only evaluate models’ performance on a problem with subproblems based on whether the model pass the whole problem.Such method will evidently loss many information since different models may have different performance on subproblems even though they all fail to solve the whole problem.
Therefore, it’s much better to locate the exact turn $t$ at which the model fails by using unit tests for each subproblem. However, directly comparing failing turns may be unfair because the topological ordering of subproblems at same depth in AST can be arbitrary. To address this issue, we propose the Pass Depth (PD) metric. For a problem with overall-depth $D$, we define its PD as follows:
$PD=\begin{cases}D-d-1,&\text{if the model makes an error on a depth-$d$ subproblem;}\\ D,&\text{if the model passes the whole problem.}\end{cases}$
We will report the Average Pass Depth (APD) by averaging the PD values across problems grouped by depth. In the case of $k$ trials, we define PD(k) as the maximum PD achieved over $k$ attempts, and we use PD(k) to compute the corresponding APD(k).
4 Experiments
4.1 Experiment Setup
In experiment part, we mainly focus on models that either demonstrate advanced general capability or are explicitly pre-trained for code generation. We evaluate both close-sourced and open-sourced models. The Closed-source models include o1-mini, GPT-4o-mini, GPT-4o, and Claude-3.5-Sonnet and the open-source models include the Qwen family (Qwen-2.5-Coder-7B-Instruct, Qwen-2.5-Coder-32B-Instruct[9], Qwen-2.5-72B-Instruct[26] and QwQ-32B[24]), the DeepSeek family (DeepSeek-v3[18] and DeepSeek-R1[6]), the Llama family (Llama-3.1-8B-Instruct and Llama-3.3-70B-Instruct[4]) and Yi-Coder-9B-Chat[20]. Together, these models cover the most widely used large-parameter and reasoning models in code generation task.
To reduce evaluation overhead and eliminate potential data leakage, we select the most recent 20% of problems in CodeFlowBench as our test set. Each model is evaluated in both multi-turn and single-turn patterns, using Pass@k and APD as the primary metrics. Detailed experimental settings is included in appendix.
4.2 Main Result
Table 1: Performances of pass@1 and Average Pass Depth of various models on CodeFlowBench. The overall average depth of problems in CodeFlowBench is 1.94.
Model pass@1 Average Pass Depth
Multi-turn Single-turn Overall Depth1 Depth2 Depth3 Depth4
Closed-Sources
o1-mini 20.8 37.8 0.541 0.233 0.581 0.818 1.125
GPT-4o-mini 13.8 22.0 0.423 0.138 0.438 0.697 1.167
GPT-4o 17.5 30.2 0.494 0.177 0.539 0.758 1.042
Claude-3.5-Sonnet 13.6 36 0.414 0.117 0.444 0.648 1.167
Open-Sources (7B-Level)
Qwen2.5-Coder-7B-Instruct 2.3 15 0.233 0.018 0.247 0.436 0.750
Llama-3.1-8B-Instruct 0.9 3.5 0.208 0.011 0.224 0.412 0.792
Yi-Coder-9B-Chat 5 13.8 0.228 0.021 0.270 0.388 0.500
Open-Sources (32B-Level)
Qwen2.5-Coder-32B-Instruct 8.6 19.8 0.316 0.067 0.342 0.570 0.917
QwQ-32B-Preview 17.3 32.7 0.487 0.261 0.539 0.570 1.042
Open-Sources (70B-Level)
Llama-3.3-70B-Instruct 15 27.6 0.448 0.163 0.465 0.733 1.042
Qwen2.5-72B-Instruct 9.1 21.3 0.301 0.110 0.314 0.497 0.583
Open-Sources (Large Model)
Deepseek-V3 18 35.7 0.529 0.219 0.549 0.836 1.208
Deepseek-R1 20.5 46.1 0.569 0.303 0.606 0.842 0.916
Performances of different LLMs
Table 1 shows the main result of our experiment. The best pass@1 achieved on CodeFlowBench is only 20.8% and the highest Average Pass Depth (APD) is also only 0.541, underscoring the benchmark is challenging enough for today’s leading LLMs. It follows our intuition that closed-source models generally outperform open-source models.However, the gap is not decisive: the open-source DeepSeek-V3 exceeds most closed-source competitors, and Llama-3.3-70B-Instruct remains highly competitive. Besides, two kinds of model consistently punch above their parameter class: specialized coding models and reasoning models. Smaller coder models like Qwen2.5-Coder and Yi-Coder deliver notably strong pass@1 scores compared to large models, while reasoning models like QwQ-32B-Preview can approach the performance of GPT-4o. These results suggest that advancing domain-specific coding architectures and refining reasoning strategies are promising directions for future performance gains.
Comparison of pass@k on multi-turn and single-turn pattern
The results in Table 1 reveal a substantial gap between the multi-turn and single-turn settings, indicating that even when facing the same underlying problem, models find multi-turn cases significantly more challenging to handle. Furthermore, we observe distinct trends in model performance across the two patterns. Models such as Claude-3.5-Sonnet and Deepseek-R1 achieve outstanding results under the single-turn setting, yet their performance in multi-turn scenarios is less remarkable. In contrast, models like GPT-4o and Llama-3.3-70B-Instruct show relatively smaller performance gaps between the two settings. Unlike the single-turn pattern, which primarily requires complex reasoning, the multi-turn pattern demands that models generate reusable and generalizable functions based on a given dependency structure and pre-implemented materials, thus requiring more comprehensive capabilities. As a result, different models exhibit varying shortcomings that need to be further addressed and enhanced.
Comparison of APD on different depth
Results from APD provide additional valuable insights. We observe that small open-source models exhibit varied strengths between low-depth problems and deep-depth problems. Models like Qwen2.5-Coder and Llama-3 perform better on problems whose depth beyond 3, whereas Yi-Coder perform better on problems whose depth is 1 or 2. In contrast, large and closed-source models with higher overall APD demonstrate more balanced performance, typically passing 20–30% depth of the whole problems in both low-depth problems and deep-depth problems. Additionally, while reasoning model like Deepseek-R1 and o1-mini achieve SOTA performance on our benchmark, it doesn’t reflect obvious advantage on deep-depth problems, indicating a possible direction for reasoning model to make thinking chain fit multi-turn cases better.
5 Discussion
While models may succeed on multi-turn tasks with a small number of turns and simple linear dependencies structure, they are nearly incapable of solving all turns correctly when the structure of problem becomes more complex.
The results in Table 1 report the average pass@1 scores across models. However, a deeper look of the problems that were successfully solved reveals a clear imbalance: most of the passed cases correspond to problems with simple dependency structures. As the problem structure becomes more complex, existing SoTA models struggle significantly and are nearly impossible for them to correctly complete all turns and solve the entire problem.
Figure 5: The pass@1 score of models on multi-turn problems with different turns.We divide models into 4 groups: closed-source models, open-source 7B-level models, open-source 32B&70B-level models and open-source large models & reasoning models.
Figure 5 presents the pass@1 scores of models on multi-turn problems across different numbers of turns. All curves, regardless of model group, exhibit a consistent trend: they begin at relatively high scores for problems with only 1 or 2 turns, but drop sharply as the number of turns increases, eventually reaching zero. For small open-source models, the scores remain low across all turn depths. Although larger open-source models and closed-source models perform significantly better on short-turn problems, their performance still degrades notably on problems with more turns. Even top-performing models such as o1-mini and Deepseek-R1, which achieve strong overall pass@1 scores, fail to solve problems with more than six turns. This result highlights a key limitation: current models still struggle to generate code in each turn that generalizes well, covers a wide range of cases, and integrates seamlessly with higher-level functions.
To further measure the complexity of dependency structure, we introduce a metric called DSC(Dependency Structure Complexity) which is defined as the ratio of its overall turns to the number of overall depth:
$\mathrm{DSC}(\text{problem})=\frac{\mathrm{OverallTurn}(\text{problem})}{\mathrm{OverallDepth}(\text{problem})}$
Recall that the overall-turn and overall-depth of a problem are derived from its AST, corresponding to the number of nodes and the depth of its AST. Based on this, we can see that a high DSC value indicates a problem with a complex dependency structure. Figure 6 presents the pass@1 scores of models across different DSC intervals. It can be observed that most models are only capable of solving problems with DSC equal to 1, which corresponds to a simple linear dependency structure. Only a few leading models, such as o1-mini, GPT-4o and Deepseek-V3, are able to solve a limited number of problems with DSC values below 1.33. All models struggle significantly when faced with problems involving more complex structures. These results highlight the current models’ limited ability to understand intricate dependency relationships and to generate functions that can effectively handle them.
Figure 6: Heatmap of models’ pass@1 scores on multi-turn problems within different DSC intervals.
Which Kind of Error Lead to LLMs’ Poor Performance in Multi-Turn Pattern?
Since a huge gap between models’ performances on multi-turn and single-turn pattern is demonstrated, we may be curious about the underlying reasons. Firstly, we define a taxonomy where errors are concluded into three types:
  • Incomplete Reasoning (IR) The model often handles only straightforward “happy-path” cases and fails to generalize. It may misinterpret or oversimplify key requirements, omit boundary or atypical cases, or choose naive algorithms whose logic or performance collapse on larger inputs. This reflects limitation of model in its reasoning ability.
  • Insufficient Globalization (IG) The function’s logic may run correctly in isolation but omits necessary imports, global constants, or shared-state interactions, preventing proper integration into the broader application or runtime. This reflects limitation of model in managing global context.
  • Instruction Misinterpretation (IM) In multi-step or multi-turn prompts, the model may solve isolated subtasks but fail to respect overall orchestration: it can overlook how to invoke provided helper functions and arrange the codeflow properly in high-level function, leading to integration failures. This reflects limitation of model in task understanding and dependency relationship handling.
Compared to the single-turn pattern, Incomplete Reasoning (IR) remains a shared challenge where the single-turn approach can leverage holistic code coordination and design to address logical issues that localized implementations may fail to anticipate; Insufficient Globalization (IG) and Instruction Misinterpretation (IM) are problems specific to the multi-turn pattern. We then conduct a study to quantify the distribution of each error type made by the LLMs. We randomly sample examples from the test set and manually annotate their error categories. The results are summarized in Table 2.
Table 2: Proportion of each error type across Deepseek-V3 and o1-mini.IR, IG and IM stand for Incomplete Reasoning, Insufficient Globalization and Instruction Misinterpretation, respectively.
Model IR(%) IG(%) IM(%) Others(%)
Deepseek-v3 47.2 14.6 32.6 5.6
o1-mini 38.7 11.8 43.0 6.5
From Table 2, we draw several insights. First, Incomplete Reasoning (IR) remains the dominant error type which is the main challenge in competition-level code generation. We need to distinguish the effect of multi-turn setting out by quantifying how much of IR can be “fixed” by switching from multi-turn pattern to single-turn pattern. Therefore, we introduce the IR Remediation Rate, which is defined as the fraction of IR errors under the multi-turn pattern that are correctly answered under single-turn evaluation. We report the IR remediation rate of two models tested above in Table 3.
Table 3: IR remediation rate of Deepseek-V3 and o1-mini.
Model IR Remediation Rate (%)
DeepSeek-v3 16.7
o1-mini 8.3
By contrast, Insufficient Globalization (IG) and Instruction Misinterpretation (IM) are largely specific to multi-turn interactions. This makes sense: CodeFlowBench tasks require multi-step orchestration, where the model must (1) decompose the high-level problem into subtasks, (2) correctly invoke pre-implemented functions and respect their dependencies, and (3) assemble an end-to-end solution that can be executed iteratively. The high rate of IM errors highlights that models still struggle with consistent instruction following across multiple turns—understandable, given that they are typically trained on isolated, single-turn examples. Finally, we observe that models often lack “global awareness” when generating low-level code: they may omit necessary imports or mishandle shared state, even though tool functions and the top-level wrapper are expected to integrate seamlessly. This underlines the importance of evaluating code generation in settings that demand both local correctness and global coherence.
Appendix B presents a more detailed case study of these errors.
6 Conclusion
In this paper, we introduce CodeFlowBench, the first benchmark specifically designed for multi-turn, iterative code generation. Comprising 5,258 complex programming problems sourced from Codeforces, CodeFlowBench is underpinned by a novel automatic data curation pipeline and a comprehensive evaluation framework that rigorously assesses a model’s ability in multi-turn and iterative code generation. Our experiment across a range of popular LLMs highlights the substantial challenges posed by this benchmark: current leading models struggle markedly with multi-turn tasks, exhibiting a pronounced performance gap compared to single-turn generation. While different models demonstrate strengths on problems with different depth, almost none can fully and correctly resolve structurally complex tasks once. Besides, we analyze key error types contributing to these shortcomings and discuss their implications. We believe CodeFlowBench not only illuminates critical limitations in existing LLMs but also provides a valuable foundation for future research aimed at closing this gap and supporting realistic development workflows.
Limitation & Future Work
While inspired by real-world development workflows, our current dataset remains confined to competition-level problems at function and class level. In future work, we plan to expand CodeFlowBench to repository-level challenges, which feature far more intricate dependency structures and demand explicit codeflow organization to ensure extensibility and maintainability. Additionally, we intend to design and train a dedicated code-generation model specifically tailored to this multi-turn, dependency-aware setting.
Appendix A Data Curation Pipeline
A.1 Detail of Problem Scraping
The original problem contain two parts. The first part is scraped from its corresponding codeforces official website. An example problem page is shown in Figure 7.
Figure 7: An example page of problems on Codeforces, which contains problem ID, title, time and memory complexity limits, problem description, input description, output description, sample tests, notes of each problem.The original problem is 196E
The second part is scraped by Problemsets.Problems [^3] https://codeforces.com/api/problemset.problems API provided by Codeforces.We record the rating and tags of each problem. Rating is a metric that reflect the difficulty of the problem and tags is a list that contains knowledge scope of the problem.
Combine the above two part, we obtain a original coding problem for CodeFlowBench, a full example is shown in Figure 8.
Figure 8: An example of original coding problem we obtained in stage I.To make the content more clear, we remove html denotions that contained in original content.The original problem is 1946E, which is used for illustration in Figure 3
A.2 Detail of Solution Scraping
During the problem-scraping process, we also collected the links under the “Contest material” section on the right side of each problem page and identified which of those led to solution blogs. Crawling these editorial pages is fairly complex, because Codeforces’ official write-ups are hosted as personal blogs whose formats can vary over time, making content extraction more challenging. Although each round’s problems (e.g. “123A,” “123B,” etc.) live on separate pages, all of the editorials for a given numeric ID usually appear on a single blog page. We therefore need to assign each sub-problem’s write-up (A, B, C, …) to its corresponding problem. To do this, our crawler first locates the distinct anchor points that mark each sub-problem section, then extracts the content between each anchor and the next as that problem’s editorial. The benefit of this approach is that, while in the end we only need the solution code itself, editorials often consist of plain text explanations, a single code snippet, or multiple code variants and languages. Relying on a purely mechanical scraper makes it difficult to isolate exactly the code we want, so it’s more effective to pass the raw editorial content to an LLM for final organization and extraction.
In general, there are two ways to locate an anchor: by URL and by text. Blogs label problems in many different forms. However, most blogs make that label into a hyperlink pointing back to the original problem, which gives us a reliable way to identify the anchor.
Therefore, our primary and most precise method is to search for a URL containing the problem ID (for example, “problemset/problem/2060/A” or “contest/blog/2060/A”) and treat its position as the anchor. Once the anchor is found, we scan the surrounding page and extract its content to obtain the editorial.Anchor scanning and recording also relies on a problem-ID reference table, which we built from the IDs of all problems scraped in the first step. Its main role is to guide the code when matching anchors: for instance, if the table shows that numeric ID 2060 has sub-IDs “A” through “G”, the scraper first reads those sub-IDs, then walks through the page using the URL-based or text-based method to record the exact anchor for each sub-ID.An example that fit URL anchor identification technique in shown in Figure 9(a).
However, there are still little parts of early editorial blogs didn’t include URL hyperlinks, so in those cases we fall back on regular-expression–based text matching wherever possible. Based on the formats we observed, we designed two main matching strategies:
  • Difficulty-label matching. A number of blogs publish all of a round’s problems on one page and mark them with labels like “Div2” or “Div1” (since most rounds contain two Div2-level problems and several Div1-level ones). To handle this, we use our problem-ID reference table to identify all sub-IDs belonging to the same numeric contest but with different difficulty levels, tag them accordingly in the table, and then, during the anchor-matching process, if the scraper detects a difficulty label it will also try to match anchors based on that label. An example is shown in Figure 9(b).
  • Problem-label matching. Beyond difficulty tags, many blogs use the literal “Problem A”, “Problem B”, etc. format. We include a specific regex pattern to detect those “Problem+sub-ID” labels and assign each section to the correct sub-problem.
Figure 9: Subfigure (a) is an example of using URL anchor identification technique. The anchor here is each subtitle displayed as "2060A-Fibonacciness". The blue font color of such subtitle indicates a URL to original problem page is setted. In practice, the existence of URL is identified by analysing the HTML code of this website.For this kind of website, we identify these subtitles as anchors and scrape the content between each subtitle.Subfigure (b) is an example of using div anchor identification technique.The anchor here is each subtitle displayed as "Div2A. Parallelepiped".These kinds of subtitle occurs in a contest round that contain two div level problems.We use reference table to project the div notation to original problem id and then identify them as corresponding anchor.
After crawling the editorial for each problem, we applied a series of processing steps to ensure quality.Firstly, we removed any editorials that were too short or empty—these problems were excluded from the dataset. To make it easier for an LLM to understand and process them, we then split each editorial into two parts: the code solution and the textual explanation, so that the model can consult the code first and then the accompanying text.The processed result is shown in Figure 10.
Figure 10: The scraped and processed solution we obtained in stage II.The original problem is 1946E
A.3 Detail of Solution Code Generation
Although we’ve already scraped the official solution for each problem, an LLM-based post-processing step is still required for two main reasons:(1) Presence of “global code segments”. Some solutions include code that isn’t encapsulated in any function. We must split the entire codebase into multiple functions and ensure that the top-level function can fully solve the problem. Since these global segments can’t be recognized during standard parsing, we rely on an LLM to reorganize the provided code so that it becomes fully parsable.(2)Early solutions exist only as text.Some of the older official solutions consist solely of textual descriptions without any runnable code. We need an LLM to convert those narratives into executable code.The prompt template for code convert is shown in Figure 11
Figure 11: The prompt template used for code convertion in stage III.The whole content of the example output data is not shown for length limitation.
The Codeforces official judging system is used to verify code correctness. We employ an automated submission bot that navigates to the Codeforces submission page [^4] https://codeforces.com/problemset/submit, fills in all required fields, and submits the solution. The site will be automatically redirects to the results page after submission, from which we scrape the verification outcome.
A.4 Detail of Subproblems Generation
While the parsing process is automatically, a LLM is still needed for generating natural language description for each subproblem.The prompt template is shown in Figure 12
Figure 12: The prompt template used for generating natural language description for each subproblem in stage III.
A example of subproblems in shown in Figure 13
Figure 13: An example of subproblems we obtained in stage IV.The solution code of 1946E contains a comb function which serve as the basic tool and is reused in solve function to address the whole problem.It’s obvious that in its AST, the comb function is the leave node in depth 1 and the solve function is the root node in depth 0.
A.5 Detail of Test Cases Generation
Overall, CodeFlowBench’s test suite is composed of two parts:
  • Top-level function tests For the final subproblem—the top-level function (e.g., main or solve) that handles overall input and output—we use the public test cases provided by the Codeforces platform.
  • Subproblem function tests For every other subproblem (i.e. functions invoked by higher-level code), we wrap each function call in a helper that redirects stdin and stdout to an internal buffer and records the resulting I/O. This can generate redundant calls for the same function, so we apply two safeguards to keep the test suite concise: deduplication of identical test cases and limiting the total number of test cases per function.These measures ensure comprehensive coverage without unnecessary duplication or excessive test-case volume.
Such method can ensure the test case generation pipeline to be automatic and the case itself to be correct.
Appendix B Error Case Study
We have defined three typical error types for test in multi-turn pattern, in this part we will introduce several example to illustrate them.We have simplify and arrange model’s outputs to make it clear to read. The content of "generated" field is model’s output. The content of "harness_result" field is the verification result by running it with test cases."1" denotes accepted, "0" denotes wrong answer and "wrong" denotes running error.
Figure 14: Example of an Incomplete Reasoning (IR) Error by Deepseek-V3.The original problem is 1799F.The algorithm produced by the model fails to find the global optimum because it greedily sorts each element according to the single-step gain—choosing either the halving operation or the subtraction operation in isolation. In reality, achieving the global optimum requires a two-dimensional dynamic-programming solution that considers both operations jointly.
Figure 15: Example 1 of an Insufficient Globalization(IG) error by Deepseek-V3.The original problem is 2005D.The model’s generated code omitted the import math statement, resulting in an error when calling gcd. This issue stems from improper handling of external imports.
Figure 16: Example 2 of an Insufficient Globalization(IG) error by Deepseek-V3.The original problem is 1734F Because the program reads new inputs and performs fresh calculations for each test case, failing to clear the cache beforehand can inadvertently reuse stale results from a previous case. This leads to incorrect output due to unintended interaction between caching and I/O.
Figure 17: Example of an Instruction Misinterpretation(IM) error by Deepseek-V3.The original problem is 1575H Although the DP routine correctly unpacks both transition and match_counts from build_transition(b), it merely treats match_counts as an ordinary list. In reality, match_counts carries two essential pieces of semantic information.(1)Match indicator.It denotes a complete match of b.(2)Backtrack hint. It’s a hint in combination with the prefix function, indicating how far the automaton should jump back after a match to continue detecting overlapping occurrences. This error is caused by models’ misinterpretation to the dependency relationship between tool function and top-level function