### v3

# for Complex Code Generation

Abstract
Modern software development demands code that is maintainable, testable, and scalable by organizing the implementation into modular components with iterative reuse of existing codes. We formalize this iterative, multi-turn paradigm as codeflow and introduce CodeFlowBench, the first benchmark designed to comprehensively evaluate LLMs’ ability to perform codeflow - implementing new functionality by reusing existing functions over multiple turns. CodeFlowBench comprises two complementary components: CodeFlowBench-Comp, a core collection of 5,000+ competitive programming problems from Codeforces updated via an automated pipeline and CodeFlowBench-Repo, which is sourced from GitHub repositories to better reflect real-world scenarios. Furthermore, a novel evaluation framework featured dual assessment protocol and structural metrics derived from dependency trees is introduced. Extensive experiments reveal significant performance degradation in multi-turn codeflow scenarios. Furthermore, our in-depth analysis illustrates that model performance inversely correlates with dependency complexity. These findings not only highlight the critical challenges for supporting real-world workflows, but also establish CodeFlowBench as an essential tool for advancing code generation research.
CodeFlowBench: A Multi-turn, Iterative Benchmark
for Complex Code Generation
Sizhe Wang1,2, Zhengren Wang1,3, Dongsheng Ma1 Yongan Yu1,4, Rui Ling1, Zhiyu Li*3, Feiyu Xiong3, Wentao Zhang*1 1Peking University 2Shanghai University of Finance and Economics 3Institute for Advanced Algorithms Research, Shanghai 4McGill University {lizy,xiongfy}@iaar.ac.cn, wentao.zhang@pku.edu.cn [Uncaptioned image] Dataset    [Uncaptioned image] Code
footnotetext: $\dagger$ Equal contribution; * Corresponding author.
The first author completed this work during an internship at Peking University.
1 Introduction
Large Language Models (LLMs) have revolutionized code generation, with benchmarks like HumanEval (chen2021evaluating) and MBPP (austin2021program) establishing foundational standards. As LLM capabilities advance, their role in real-world software development has expanded beyond solving toy problems to supporting complex workflows (jiang2024survey; jin2024llms; liu2024large). Modern benchmarks such as SWE-Bench (jimenez2023swe; li2024devbench) now emphasize practical scenarios like bug fixing. However, current benchmarks (gu2024cruxeval; huang2024effibench; wang2025maintaincoder) still overlook the critical aspect of real-world development: the multi-turn and iterative codeflow scenario.
The CodeFlow Task
In modern software engineering, multi-turn and iterative workflows are becoming increasingly prevalent, as the cornerstones of best practices like agile development (larman2004agile; abrahamsson2017agile; cram2019agile). By breaking down complex tasks into manageable subproblems, progressively refine solutions and reuse modular functions, developers can achieve faster delivery, reduced redundancy and enhanced maintainability in teamwork (haefliger2008code; feitosa2020code). For example, React’s core package alone sees over 37 million weekly downloads across 2,300+ dependent modules, illustrating the productivity gains of modular reuse boduch2019react. As shown in Fig. 1 and 2, by building solutions via traversing the function dependency tree from the bottom up, the structured codeflow can both enables parallel development and improves readability, testability and maintainability through modular and responsibility boundaries. To integrate effectively into codeflows, LLMs must learn to “look before and after”, namely accurately leveraging pre-implemented functions while ensuring the modularity of codes to generate.
Figure 1: A currency arbitrage example contrasting the modular codeflow workflow (top) with a monolithic single-turn implementation (bottom).
Motivation
Despite the growing demands, current benchmarks have not fully captured the multi-turn and iterative aspects of codeflow. Firstly, most benchmarks, such as HumanEval and MBPP, focus only on single-turn code generation. While recent benchmarks like BigCodeBench (zhuo2024bigcodebench) and SWE-Bench (jimenez2023swe) have begun to incorporate practical development scenarios, they still remain at the level of single-turn code modification (or two-turn code generation), leaving the multi-turn generation capabilities unknown. Secondly, the few existing multi-turn benchmarks such as MTPB (nijkamp2022codegen), only focus on single-function programming, lacking both unit tests and sufficient complexity to reflect real-world dependencies. Finally, due to the absence of an update mechanism, previous static datasets risk contamination and unreliable assessment. Therefore, there is a pressing need for a more challenging, well-annotated and frequently updated benchmark specially designed for codeflow.
To bridge this gap, we introduce CodeFlowBench, a benchmark suite that integrates two complementary datasets: CodeFlowBench-Comp, which is sourced from competitive programming platforms like Codeforces [^1] https://codeforces.com/ and CodeFlowBench-Repo, which is derived from diverse real-world GitHub repositories via DomainEval (zhu2025domaineval). Utilizing a unified automated pipeline, CodeFlowBench transforms raw data into sequences of multi-turn, iterative subproblems, accompanied by verified solutions and test cases. This dual-component architecture ensures a broad evaluative scope, spanning reasoning-intensive algorithmic challenge and labor-intensive software engineering. Specifically, CodeFlowBench-Comp provides high-difficulty, high-quality problems with periodic updates, while CodeFlowBench-Repo injects real-world complexity into the assessment. To further reveal model deficiencies, CodeFlowBench introduces a specialized evaluation framework and incorporates structural labels and metrics derived from dependency trees, enabling a nuanced analysis of multi-turn performance.
Contributions
Our contributions are threefold:
  • Pipeline Innovation: We developed a universal data curation pipeline grounded in function dependency analysis. This fully-automated and lightweight framework decomposes monolithic solutions into multi-turn, iterative coding problems that necessitate strategic code reuse. Its design is rigorously verifiable and highly extensible, enabling the transformation of diverse source code into structured codeflow-style challenges.
  • Benchmark Construction: We introduce CodeFlowBench, the first benchmark to evaluate the iterative, multi-turn code generation capabilities of LLMs. By integrating two complementary datasets: CodeFlowBench-Comp (algorithmic depth) and CodeFlowBench-Repo (real-world domain breadth), we provide a comprehensive testbed to examine software engineering proficiency.
  • Evaluation Design & Insights: We propose a novel evaluation framework that directly contrasts multi-turn and single-turn generation patterns. By introducing structural metrics, such as Average Pass Depth (APD) and Dependency Structure Complexity (DSC), derived from dependency trees, our framework captures the unique complexities of iterative tasks. Extensive experiments reveal a significant performance degradation in codeflow scenarios, even among state-of-the-art reasoning models, highlighting a critical frontier for future research and the need for more advanced codeflow programming capabilities.
2 Related Work
Code Generation Benchmarks
The landscape of code generation benchmarks has evolved from simpler to more complex tasks to keep pace with the rapid development of LLMs (li2022competition; quan2025codeelo; yu2024humaneval), but still fail to comprehensively capture the multi-turn and iterative features of real-world scenarios. Early works like HumanEval (chen2021evaluating) and MBPP (austin2021program) focus on standalone functions with low complexity and limited dependency environments. Recent benchmarks have emerged to evaluate more complex and realistic scenarios, yet have obvious limitations: most benchmarks such as APPS (hendrycks2021measuring), LiveCodeBench (jain2024livecodebench) and SWE-Bench (jimenez2023swe) are limited to single-turn code generation or modification. For the few existing multi‑turn benchmarks, MTPB (nijkamp2022codegen) focus on overly simplistic single-function programming without paired unit tests, while InterCode (yang2023intercode) discusses interactive coding with execution feedback. In stark contrast, codeflow structures the development into multi-turn processes, ensuring each component is maintainable, testable and reusable. These limitations highlight the crucial gaps for codeflow benchmarking, and the CodeFlowBench pioneers this research line.
Code Generation LLMs
Recent years have witnessed unprecedented progress in code generation capabilities of LLMs. Early works such as Codex (chen2021evaluating) and AlphaCode (li2022competition) demonstrated proficiency in tasks ranging from code completion to competition-level problem solving. With the scaling up of pre‑trained models, exemplified by GPT (gpt5), Claude (claude4), Deepseek‑Coder (guo2024deepseek) and Qwen‑Coder (yang2025qwen3), these advanced models have impressive performance across various programming tasks, languages and domains. Building on these foundations, the code generation capabilities have further advanced through instruction tuning and agent frameworks. Models such as WizardCoder (luo2023wizardcoder) and Magicoder (wei2023magicoder) leverage instruction tuning to improve intent alignment and interactive dialogue capabilities, while agent frameworks like AgentCoder (huang2023agentcoder) and MapCoder (islam2024mapcoder) enable autonomous planning, iterative refinement, and self-evaluation. Despite advancements, the community still remains unknown about "how well and how deeply LLMs can perform codeflow"—a critical paradigm for real-world software engineering. Our CodeFlowBench thus provides a principled framework for advancing both model development and evaluation.
3 CodeFlowBench
In this section, we introduce CodeFlowBench’s data curation pipeline and evaluation framework. Fig. 2 provides an example of CodeFlowBench.
Figure 2: An illustrative example from CodeFlowBench (Source: Codeforces 2055E). The figure illustrates the problem definition with its dependency structure (top), the iterative subproblem decomposition (middle), and the corresponding LLM interaction workflow (bottom).
3.1 Data Curation Pipeline
We designed an automated, lightweight data-curation pipeline to generate complex multi-turn coding problems. As shown in Fig. 3, the pipeline mainly consists a data preparation phase and a subproblem generation phase.
Figure 3: The data curation pipeline of CodeFlowBench. In Data Preparation Phase (Steps I–III), we scrape problem statements and editorial information, followed by generating and validating the canonical solution codes. In Subproblem Generation Phase (Steps IV–VI), we decompose solutions into subproblems via AST parsing, generate corresponding test cases, and perform automatic complexity labeling to achieve the final CodeFlowBench.
Data Preparation Phase
To address the lack of reasoning-intensive programming datasets tailored for codeflow scenarios, this phase collects programming challenges from Codeforces to construct CodeFlowBench-Comp. We first scrape problem metadata and official editorial content from the platform. To ensure data consistency, raw editorials are processed using DeepSeek-R1 (guo2025deepseek) to generate modular, parsable solution code. These generated solutions undergo rigorous verification via the Codeforces submission system; only implementations that pass all official test cases are retained. By anchoring our dataset in official editorials and validating them against a live judge, we guarantee high-quality, correct codeflow implementations for every problem. Further specifics are provided in Appendix A.1A.3.
Subproblem Generation Phase
This phase functions as a general framework for transforming monolithic solutions into the codeflow format. First, we parse the AST of the verified solution to extract and topologically sort function dependencies. We treat each function as a distinct subproblem, employing Deepseek-V3 (liu2024deepseek) to back-translate its logic into a natural language statement. Next, we execute the full solution against public test cases, instrumenting function calls to capture input-output pairs as ground-truth unit tests for each subproblem. Finally, we annotate problem complexity such as dependency structure complexity, metrics that correlate strongly with official difficulty ratings. Implementation details are provided in Appendix A.4A.5.
Remark. To ensure our pipeline is not biased by the LLMs used for subproblem description and code generation, we conducted several ablation studies. We assessed description consistency across models and validated quality via Human-LLM cross-validation. Furthermore, we verified the structural stability of code decomposition across different LLMs. Ablation details are found in Appendix A.6, and cost estimates for pipeline running are in Appendix A.7 for interested readers.
3.2 Benchmark Details
CodeFlowBench-Comp
This component comprises 5,258 competitive programming problems designed to test algorithmic reasoning depth. It emphasizes multi-step logic with up to $\geq$7 interaction turns. Similarly, dependency structures are non-trivial, concentrated at depth 2, indicating that nested function calls are the norm rather than the exception. To validate the difficulty of these structures, we correlated our metrics with official Codeforces ratings. As illustrated in Fig. 12 (Appendix B for space reason), there is a strong positive correlation: as interaction turns and dependency depths increase, the average problem rating rises sharply, approaching 2,900 (Grandmaster level) for tasks with $\geq$7 turns. These statistics confirm that our structural metrics effectively capture the intrinsic complexity of logical problems.
CodeFlowBench-Repo
Complementing the algorithmic focus, CodeFlowBench-Repo evaluates software engineering proficiency in real-world contexts. It is constructed by applying our Subproblem Generation pipeline to DomainEval (zhu2025domaineval), utilizing high-quality GitHub repositories (>>100 stars) as the source. Our curated dataset spans five specialized domains: Basic, Cryptography, Network, System, and Visualization. Through our pipeline, we extended these original contexts into complex codeflow scenarios that reach up to 9 interaction turns and a dependency depth of 4. The final distribution highlights significant structural complexity: 56.6% of problems require multi-turn iterative generation, forcing models to maintain reasoning continuity, while 54.3% involve deep dependency chains, requiring the model to correctly orchestrate hierarchical function calls. Detailed construction protocols and statistics are provided in Appendix C.1, C.2 and C.3.
3.3 Evaluation Framework
Task Definition
Our evaluation follows best practices defined in HumanEval (chen2021evaluating), where models are required to implement a given function in each round. However, CodeFlowBench introduces distinct differences in the supporting materials provided to the models. We include the function signature $F_{t}$, problem description $S_{t}$ and pre-implemented functions for code reuse. Additionally, we provide an overarching background description of the problem as $B$. This reflects the real-world scenario where developers often possess overall understanding of the entire problem when implementing code incrementally. As illustrated in Fig. 2, models must implement target functions, $C_{t}$ at turn $t$, by leveraging pre-implemented components from prior turns $\{C_{1},...,C_{t-1}\}$.
For baseline comparison, we define the single-turn variant where the model generates all components simultaneously. The mathematical expressions of both settings are presented below:
$C_{1:T}$ $\leftarrow p\Bigl(\cdot\mid F_{1:T},S_{1:T},B\Bigr).$
$C_{t}$ $\leftarrow p\Bigl(\cdot\mid F_{t},S_{t},C_{<t},B\Bigr),\quad 1\leq t\leq T.$
Performance Metrics
We adopt widely-used Pass@k (chen2021evaluating) as the main metric for both multi-turn and single-turn cases, but augment it with novel diagnostics for multi-turn analysis. While Pass@k evaluates final success rates, the coarse-grained nature fails to capture partial progress in failed attempts -two models may fail solve the entire problem at different stages but receive identical scores.
To address this limitation, we propose a new metric, Pass Turn ($PT$), which identifies the exact turn at which a model fails by leveraging unit tests for each subproblem. However, only considering failing turns may be biased due to the arbitrary topological ordering at the same depth. We further define, Pass Depth (PD=DdPD=D-d), for bottom-up programming, where $d$ and $D$ are working depth and total depth. For statistical significance, we define the Average Pass Turn (APT) and Average Pass Depth (APD) by averaging across problems grouped by turn or depth. For multiple trials, we define APD@k and APT@k following the Pass@k expression. These metrics clearly distinguishing a model’s strong performance on shallow steps from its potential struggles with deep dependencies. Please refer to Appendix D for more details.
4 Experiments
Model Pass@1 (%) APD@1 (Average Pass Depth)
Multi-Turn Single-Turn Overall Depth1 Depth2 Depth3 Depth4
CodeFlowBench-Comp
Closed-Source
GPT-5 26.5 37.6 0.762 0.185 0.866 1.306 1.542
GPT-5-mini 19.1 38.0 0.655 0.103 0.748 1.200 1.375
o3-mini 22.7 38.9 0.570 0.322 0.585 0.818 1.250
o1-mini 20.8 37.8 0.541 0.233 0.581 0.818 1.125
GPT-4.1-mini 24.4 38.7 0.602 0.265 0.673 0.873 1.042
GPT-4o-mini 13.8 22.0 0.423 0.138 0.438 0.697 1.167
Gemini-2.5-flash-thinking 22.7 39.8 0.642 0.167 0.761 0.981 1.208
Gemini-3-flash-thinking 48.4 65.5 1.120 0.445 1.303 1.625 1.625
Claude-4-Sonnet 27.3 42.9 0.830 0.185 0.893 1.250 1.375
Claude-4.5-Sonnet 31.6 47.9 0.857 0.222 1.008 1.377 1.521
Open-Source (7B)
Llama-3.1-8B-Instruct 0.9 3.5 0.208 0.011 0.224 0.412 0.792
Qwen2.5-Coder-7B-Instruct 2.3 15.0 0.233 0.018 0.247 0.436 0.750
Qwen3-8B 9.1 18.3 0.406 0.088 0.444 0.738 1.125
Open-Source (32B)
Qwen2.5-Coder-32B-Instruct 8.6 19.8 0.316 0.067 0.342 0.570 0.917
Qwen3-Coder-30B-A3B 11.6 26.0 0.405 0.100 0.458 0.681 0.917
Open-Source (70B)
Llama-3.3-70B-Instruct 15.0 27.6 0.448 0.163 0.465 0.733 1.042
Qwen2.5-72B-Instruct 9.1 21.3 0.301 0.110 0.314 0.497 0.583
Open-Source (Large)
Deepseek-V3 18.0 35.7 0.529 0.219 0.549 0.836 1.208
Deepseek-R1 20.5 46.1 0.569 0.303 0.606 0.842 0.916
CodeFlowBench-Repo
Closed-Source
GPT-5 34.7 48.8 0.898 0.483 1.020 1.385 3.200
GPT-5-mini 28.1 53.5 0.740 0.362 0.725 1.615 3.000
o3-mini 28.3 51.2 0.850 0.466 0.902 1.615 2.800
o1-mini 21.4 48.8 0.817 0.362 0.860 1.615 3.600
GPT-4.1-mini 27.6 48.0 0.724 0.431 0.705 1.692 1.800
GPT-4o-mini 24.6 43.3 0.667 0.414 0.800 1.000 1.400
Gemini-2.5-flash-thinking 26.1 44.6 0.765 0.375 0.833 1.167 3.600
Gemini-3-flash-thinking 33.9 56.7 0.905 0.500 1.020 1.308 3.400
Claude-4-Sonnet 33.9 57.6 0.850 0.461 0.980 1.385 2.600
Claude-4.5-Sonnet 33.1 56.7 0.913 0.448 1.039 1.615 3.200
Open-Source (7B)
Llama-3.1-8B-Instruct 17.3 39.6 0.417 0.310 0.451 0.615 0.800
Qwen2.5-Coder-7B-Instruct 19.7 39.1 0.528 0.310 0.647 0.692 1.400
Qwen3-8B 18.1 36.2 0.394 0.293 0.450 0.462 0.800
Open-Source (32B)
Qwen2.5-Coder-32B-Instruct 25.2 45.7 0.646 0.379 0.784 1.000 1.400
Qwen3-Coder-30B-A3B 23.6 42.5 0.543 0.414 0.470 1.385 0.600
Open-Source (70B)
Llama-3.3-70B-Instruct 22.1 39.6 0.614 0.345 0.522 1.846 1.400
Qwen2.5-72B-Instruct 25.2 43.3 0.716 0.397 0.745 1.230 2.800
Open-Source (Large)
Deepseek-V3 29.1 45.1 0.771 0.483 0.745 1.538 2.400
Deepseek-R1 29.4 52.0 0.770 0.431 0.760 1.538 2.800
Table 1: Performance on CodeFlowBench-Comp and CodeFlowBench-Repo across different model scales. Both datasets report Pass@1 and APD@1 (Average Pass Depth) across varying depths.
Experiment Setup
For comprehensive experiments, we evaluate both close-sourced and open-sourced models. The closed‑source models include the GPT (gpt5), Gemini (gemini3) and Claude families (claude4). The open‑source models include the Qwen (yang2025qwen3), DeepSeek (liu2024deepseek; guo2025deepseek) and Llama families (grattafiori2024llama). We experiments on CodeFlowBench-Comp and CodeFlowBench-Repo under identical settings. For CodeFlowBench-Comp, we selected a test subset comprising the most recent 1,000 problems to eliminate potential data leakage and decrease evaluation overhead while maintaining statistical significance. All models are evaluated in both multi-turn and single-turn scenarios, utilizing Pass@k and APD@k as the primary metrics. Please refer to Appendix E.1 and E.2 for implementation details.
4.1 Main Experiments
Table 1 presents the main experiment results. For the majority of LLMs, the Pass@1 in the multi-turn scenario remains below 30%, underscoring CodeFlowBench’s difficulty. The performance across CodeFlowBench also reveals divergent model specializations. For example, Gemini-3-flash performs best on CodeFlowBench-Comp while GPT-5 outperforms on CodeFlowBench-Repo. Furthermore, we present more analysis as follows.
Multi-Turn versus Single-Turn
Table 1 reveals a substantial performance gap between multi-turn and single-turn scenarios across both CodeFlowBench-Comp and CodeFlowBench-Repo. This consistent degradation underscores the inherent complexity of multi-turn generation, which necessitates long-context coherence and robust dependency management. Notably, we observe a divergence in model behaviors: while reasoning-enhanced models like Deepseek-R1 and o3-mini achieve impressive single-turn scores, they suffer performance drops exceeding 50% in multi-turn settings. In contrast, models such as GPT-5 demonstrate greater stability with narrower gaps. This distinction suggests that single-turn metrics largely reflect peak reasoning capabilities, whereas multi-turn evaluation exposes previously under-explored limitations in iterative development and dependency management.
Model CodeFlowBench-Comp CodeFlowBench-Repo
Overall Turn 1 Turn 2 Turn 3 Turn 4 Overall Turn 1 Turn 2 Turn 3+
Closed-Source
GPT-5 0.832 0.185 0.951 1.224 1.781 0.890 0.509 1.049 1.000
GPT-5-mini 0.717 0.103 0.832 1.158 1.345 0.520 0.363 0.537 0.667
o3-mini 0.600 0.322 0.632 0.777 0.983 0.780 0.455 0.854 1.056
o1-mini 0.581 0.233 0.645 0.798 0.879 0.675 0.345 0.725 0.944
GPT-4.1-mini 0.646 0.265 0.760 0.803 1.034 0.701 0.436 0.732 1.056
GPT-4o-mini 0.467 0.137 0.501 0.638 0.982 0.571 0.418 0.650 0.778
Gemini-2.5-flash-thinking 0.695 0.167 0.855 0.923 1.218 0.687 0.364 0.879 0.600
Gemini-3-flash-thinking 1.178 0.445 1.386 1.617 1.800 0.905 0.491 1.000 1.000
Claude-4-Sonnet 0.830 0.185 0.990 1.153 1.564 0.866 0.455 1.024 1.389
Claude-4.5-Sonnet 0.927 0.222 1.104 1.291 1.704 0.961 0.418 1.146 1.167
Open-Source (7B)
Llama-3.1-8B-Instruct 0.232 0.011 0.245 0.404 0.534 0.370 0.327 0.439 0.389
Qwen2.5-Coder-7B-Instruct 0.258 0.018 0.270 0.394 0.638 0.528 0.327 0.512 0.778
Qwen3-8B 0.452 0.075 0.490 0.692 1.036 0.362 0.309 0.488 0.333
Open-Source (32B)
Qwen2.5-Coder-32B-Instruct 0.352 0.067 0.391 0.532 0.569 0.583 0.364 0.780 0.833
Qwen3-Coder-30B-A3B 0.449 0.100 0.508 0.699 0.836 0.496 0.418 0.561 0.333
Open-Source (70B)
Llama-3.3-70B-Instruct 0.493 0.163 0.515 0.681 1.000 0.516 0.345 0.553 0.750
Qwen2.5-72B-Instruct 0.330 0.110 0.350 0.452 0.517 0.669 0.364 0.780 0.889
Open-Source (Large)
Deepseek-V3 0.572 0.219 0.622 0.750 0.966 0.709 0.455 0.829 0.722
Deepseek-R1 0.609 0.304 0.677 0.766 0.966 0.675 0.418 0.925 0.611
Table 2: Pass Turn comparison across models on CodeFlowBench-Comp and CodeFlowBench-Repo. Consistent with Table 1, the overall performance distribution shows that models with high APD@1 scores also exhibit high APT@1 scores, validating the precision of our metric.
Depth-Wise Performance
Our analysis across dependency depths reveals a dichotomy in scaling laws. First, in shallow dependencies, reasoning paradigms supersede parameter scale. Efficient reasoning models, such as o3-mini, significantly outperform larger base models like GPT-4o on CodeFlowBench-Comp, suggesting that logical deduction is more critical than raw capacity for immediate or localized tasks. Second, deep dependencies function as a capacity filter, where model scale becomes decisive. In CodeFlowBench-Repo, Qwen2.5-72B-Instruct doubles the performance of its 7B and 32B counterparts, indicating that maintaining coherence across extended dependency chains requires substantial parameter capacity. Finally, SOTA models like Gemini-3-flash-thinking and GPT-5 demonstrate full-spectrum dominance, eliminating trade-offs by combining agile reasoning for local logic with the extensive capacity required for global integration.
From another perspective, we report APT@1 in Table 2. Notably, the APT trends closely align with APD, reinforcing our observations of models’ depth-wise performance.
4.2 Analysis and Discussion
Dependency Structure Challenges in Multi-turn Scenarios
A deeper analysis of solved problems reveals a striking imbalance: the majority of correctly addressed cases correspond to problems with simple, linear dependency structures (e.g., shallow call graphs or sequential compositions). However, as problem architectures evolve toward modular and hierarchical dependencies (e.g., nested function calls, interdependent components), models exhibit significant performance degradation. This phenomenon is empirically validated in Figure 4, which illustrates the Pass@1 scores across varying turn counts. The consistent performance trajectory demonstrates the inherent challenges in multi-turn code generation, where curves initially high for 1-2 turn problems and followed by sharp decline as turn counts increase. Even top-performing models fail to solve problems requiring more than six turns. This underscores the critical limitation to balance local correctness and global integration across iterative development cycles.
Figure 4: Models’ Pass@1 results on multi-turn problems grouped by model categories and turn number.
To quantity dependency structure complexity, we introduce the Dependency Structure Complexity (DSC) metric, defined as the ratio of total turns to the maximum depth in the AST. Figure 16 presents the models’ performance across different DSC intervals, revealing that most models perform well on problems with linear dependency structures but struggle significantly as the dependency structure becomes more complex. For space reason, we leave more discussion in Appendix F.1.
Fine-Grained Error Types in Multi-turn Generation
Given the significant performance gap between models in multi-turn and single-turn scenarios, we conducted studies to identify the underlying reasons. We categorized errors into three primary types: (1) Incomplete Reasoning (IR): Models often handle only straightforward "happy-path" cases and fail to generalize. They may oversimplify key requirements, omit boundary or atypical cases, or choose naive algorithms whose logic or performance collapses on larger inputs. This reflects a limitation in the models’ reasoning abilities. (2) Insufficient Globalization (IG): While a function’s logic may run correctly in isolation, it may omit necessary imports, global constants, or shared-state interactions, preventing proper integration into the broader application or runtime. This indicates a limitation in the models’ ability to manage global context. (3) Instruction Misinterpretation (IM): Given multi-turn prompts, models could solve isolated subproblems but lack a coherent understanding of the overarching goal. Typical failures include misusing helper functions or implementing disorganized code within the top-level function, i.e. incorrect integration of components. To quantify the distribution of errors, we randomly sampled examples from CodeFlowBench-Comp and CodeFlowBench-Repo. The error categories are manually annotated by authors. The proportion is calculated in Table 3.
Model Dataset IR(%) IG(%) IM(%) Others(%)
DeepSeek-V3 Comp 47.2 14.6 32.6 5.6
Repo 32.1 22.6 44.0 1.2
o1-mini Comp 38.7 11.8 43.0 6.5
Repo 23.7 18.3 55.9 2.2
Gemini-3-Flash Comp 52.3 20.5 27.3 0.0
Repo 34.6 21.8 41.0 2.6
Table 3: Error distribution across models and benchmarks. Values indicate the percentage of each error category relative to the total number of errors identified in the sampled dataset for each model.
Table 3 offers several insights. First, the distinct error distributions confirm that our benchmark components target divergent primary challenges: CodeFlowBench-Comp is dominated by IR and CodeFlowBench-Repo emphasizes global context which shifting the burden to IG and IM. Second, distinct bottlenecks constrain models of different tiers. DeepSeek-V3 and o1-mini are primarily limited by IM, suggesting their performance is capped at instruction comprehension and adherence. In contrast, for SOTA models like Gemini-3-Flash, the bottleneck shifts back to IR, indicating task understanding becomes minor issues. Detailed case studies are provided in Appendix F.2.
5 Conclusion
This paper introduces CodeFlowBench, the first benchmark specifically designed to evaluate multi-turn iterative code generation capabilities in realistic development workflows, i.e., the codeflow. CodeFlowBench consists of two subsets: CodeFlowBench-Comp for multi-step algorithmic reasoning via competition challenges, and CodeFlowBench-Repo for evaluating generalization in real-world software contexts. Our benchmark makes three key contributions: (1) an automated pipeline for decomposing complex problems into dependency-aware subproblems with paired unit tests, (2) a novel evaluation framework with proposed structural metrics, such as APD@k and DSC, to quantify multi-turn performance, and (3) the discovery of substantial performance gaps between multi-turn and single-turn scenarios (up to 50% performance degradation). Our fine-grained analysis identifies dominant failure modes and provides insights for further advancements. Extensive experiments across 19 popular LLMs highlight the substantial challenges posed by both the codeflow task and our benchmark. We believe CodeFlowBench not only illuminates critical limitations in existing LLMs but also paves the way for more realistic and powerful code generation systems.
Limitations
We acknowledge two limitations that present opportunities for future expansion. First, while the pipeline’s subproblem generation phase is general, the data preparation phase is currently specific. To achieve universal compatibility, we can consider implementing an agentic framework for automatically adapting diverse codebases. This will expand our pipeline’s coverage to a broader range of data sources beyond Codeforces and GitHub, such as Bitbucket, establishing a truly source-agnostic, end-to-end framework. Second, we currently lack models specifically fine-tuned for the codeflow task. Future work will consider fine-tuning specialized models and agentic systems on this data to assist developers in real-world, iterative programming scenarios. We also aim to extend our codeflow paradigm to include iterative debugging loops. By training specialized systems on these dynamic scenarios, we seek to enable robust and sustainable code implementation and maintenance, allowing models to effectively resolve requirements through multi-turn refinement.
Ethical Consideration
Codeforces provides an official “Codeforces materials usage license” [^2] https://codeforces.com/blog/entry/967. We have taken great care to make our pipeline fully comply with all licensing requirements and community guidelines:
Firstly, all problems we use are officially published by Codeforces, and we provide clear attribution to Codeforces. Our usage is strictly non-commercial and academic, in accordance with the license’s permitted use cases. Our data pipeline strictly avoids scraping any user-submitted code, forum discussion content or third-party solutions. We only collect officially published materials, namely problem statements and editorials.
Secondly, officially published Codeforces materials have long been used for widely recognized code generation benchmarks such as CodeContest (li2022competition), LiveCodeBench (jain2024livecodebench) and CodeElo (quan2025codeelo). Our work follows this community tradition and with the utmost respect for Codeforces.
Besides, we will also make our code properly hosted and clearly state its intended use to ensure responsible data collection.
Appendix
Appendix A Data Curation Pipeline
A.1 Detail of Problem Scraping
The original problem contain two parts. The first part is scraped from its corresponding codeforces official website. An example problem page is shown in Figure 5.
Figure 5: An example page of problems on Codeforces, which contains problem ID, title, time and memory complexity limits, problem description, input description, output description, sample tests, notes of each problem.The original problem is 196E
The second part is scraped by Problemsets.Problems [^3] https://codeforces.com/api/problemset.problems API provided by Codeforces.We record the rating and tags of each problem. Rating is a metric that reflect the difficulty of the problem and tags is a list that contains knowledge scope of the problem.
Combine the above two part, we obtain a original coding problem for CodeFlowBench, a full example is shown in Figure 6.
Figure 6: An example of original coding problem we obtained in stage I.To make the content more clear, we remove html denotions that contained in original content. The original problem is 1946E, which is used for illustration in Figure 3
A.2 Detail of Solution Scraping
During the problem‐scraping process, we also collected the links under the “Contest material” section on the right side of each problem page and identified which of those led to solution blogs. Crawling these editorial pages is fairly complex, because Codeforces’ official write‐ups are hosted as personal blogs whose formats can vary over time, making content extraction more challenging. Although each round’s problems (e.g. “123A,” “123B,” etc.) live on separate pages, all of the editorials for a given numeric ID usually appear on a single blog page. We therefore need to assign each sub‐problem’s write‐up (A, B, C, …) to its corresponding problem. To do this, our crawler first locates the distinct anchor points that mark each sub‐problem section, then extracts the content between each anchor and the next as that problem’s editorial. The benefit of this approach is that, while in the end we only need the solution code itself, editorials often consist of plain text explanations, a single code snippet, or multiple code variants and languages. Relying on a purely mechanical scraper makes it difficult to isolate exactly the code we want, so it’s more effective to pass the raw editorial content to an LLM for final organization and extraction.
In general, there are two ways to locate an anchor: by URL and by text. Blogs label problems in many different forms. However, most blogs make that label into a hyperlink pointing back to the original problem, which gives us a reliable way to identify the anchor.
Therefore, our primary and most precise method is to search for a URL containing the problem ID (for example, “problemset/problem/2060/A” or “contest/blog/2060/A”) and treat its position as the anchor. Once the anchor is found, we scan the surrounding page and extract its content to obtain the editorial.Anchor scanning and recording also relies on a problem‑ID reference table, which we built from the IDs of all problems scraped in the first step. Its main role is to guide the code when matching anchors: for instance, if the table shows that numeric ID 2060 has sub‑IDs “A” through “G”, the scraper first reads those sub‑IDs, then walks through the page using the URL‑based or text‑based method to record the exact anchor for each sub‑ID.An example that fit URL anchor identification technique in shown in Figure 7(a).
However, there are still little parts of early editorial blogs didn’t include URL hyperlinks, so in those cases we fall back on regular‑expression–based text matching wherever possible. Based on the formats we observed, we designed two main matching strategies:
  • Difficulty‑label matching. A number of blogs publish all of a round’s problems on one page and mark them with labels like “Div2” or “Div1” (since most rounds contain two Div2‑level problems and several Div1‑level ones). To handle this, we use our problem‑ID reference table to identify all sub‑IDs belonging to the same numeric contest but with different difficulty levels, tag them accordingly in the table, and then, during the anchor‑matching process, if the scraper detects a difficulty label it will also try to match anchors based on that label. An example is shown in Figure 7(b).
  • Problem‑label matching. Beyond difficulty tags, many blogs use the literal “Problem A”, “Problem B”, etc. format. We include a specific regex pattern to detect those “Problem+sub‑ID” labels and assign each section to the correct sub‑problem.
Figure 7: Subfigure (a) is an example of using URL anchor identification technique. The anchor here is each subtitle displayed as "2060A-Fibonacciness". The blue font color of such subtitle indicates a URL to original problem page is setted. In practice, the existence of URL is identified by analysing the HTML code of this website.For this kind of website, we identify these subtitles as anchors and scrape the content between each subtitle.Subfigure (b) is an example of using div anchor identification technique.The anchor here is each subtitle displayed as "Div2A. Parallelepiped".These kinds of subtitle occurs in a contest round that contain two div level problems.We use reference table to project the div notation to original problem id and then identify them as corresponding anchor.
After crawling the editorial for each problem, we applied a series of processing steps to ensure quality.Firstly, we removed any editorials that were too short or empty—these problems were excluded from the dataset. To make it easier for an LLM to understand and process them, we then split each editorial into two parts: the code solution and the textual explanation, so that the model can consult the code first and then the accompanying text.The processed result is shown in Figure 8.
Figure 8: The scraped and processed solution we obtained in stage II. The original problem is 1946E
A.3 Detail of Solution Code Generation
Although we’ve already scraped the official solution for each problem, an LLM-based post-processing step is still required for two main reasons:(1) Presence of “global code segments”. Some solutions include code that isn’t encapsulated in any function. We must split the entire codebase into multiple functions and ensure that the top-level function can fully solve the problem. Since these global segments can’t be recognized during standard parsing, we rely on an LLM to reorganize the provided code so that it becomes fully parsable.(2)Early solutions exist only as text.Some of the older official solutions consist solely of textual descriptions without any runnable code. We need an LLM to convert those narratives into executable code.The prompt template for code convert is shown in Figure 9
Figure 9: The prompt template used for code convertion in stage III.The whole content of the example output data is not shown for length limitation.
The Codeforces official judging system is used to verify code correctness. We employ an automated submission bot that navigates to the Codeforces submission page [^4] https://codeforces.com/problemset/submit, fills in all required fields, and submits the solution. The site will be automatically redirects to the results page after submission, from which we scrape the verification outcome.
A.4 Detail of Subproblems Generation
While the parsing process is automatically, a LLM is still needed for generating natural language description for each subproblem.The prompt template is shown in Figure 10
Figure 10: The prompt template used for generating natural language description for each subproblem in stage III.
A example of subproblems in shown in Figure 11.
Figure 11: An example of subproblems we obtained in stage IV.The solution code of 1946E contains a comb function which serve as the basic tool and is reused in solve function to address the whole problem.It’s obvious that in its AST, the comb function is the leave node in depth 1 and the solve function is the root node in depth 0.
A.5 Detail of Test Cases Generation
Overall, CodeFlowBench’s test suite is composed of two parts:
  • Top-level function tests For the final subproblem—the top-level function (e.g., main or solve) that handles overall input and output—we use the public test cases provided by the Codeforces platform.
  • Subproblem function tests For every other subproblem (i.e. functions invoked by higher-level code), we wrap each function call in a helper that redirects stdin and stdout to an internal buffer and records the resulting I/O. This can generate redundant calls for the same function, so we apply two safeguards to keep the test suite concise: deduplication of identical test cases and limiting the total number of test cases per function.These measures ensure comprehensive coverage without unnecessary duplication or excessive test-case volume.
Such method can ensure the test case generation pipeline to be automatic and the case itself to be correct.
A.6 Ablation Study of LLM usage
First, to check its influence on the content and style of Subproblem Descriptions, we compared Deepseek-V3 against GPT-4o-mini, Gemini-2.0-Flash and o3-mini. As shown in Table 4, the high semantic overlap and consistent stylistic metrics indicate that descriptions are primarily driven by the underlying code logic rather than model-specific habits. The main content of subproblems is not highly influenced by model choice.
Second, to check its influence on the structure of LLM-generated solution code, we re-generated solutions using o3-mini, Claude-3.7-Sonnet, and Gemini-2.5-Pro. The resulting complexity metrics (Table 5) were identical across all models, confirming that the model produces almost identical decomposition results, indicating very minor bias between different LLMs.
Finally, to verify that our subproblems are natural and representative of human problem-solving, we conducted a Human-LLM cross-validation by constructing a creteria including five aspects. As shown in Table 6, both human and model judges consistently rated the subproblems as high-quality, which confirms that our automated pipeline produces professional, human-aligned problem decompositions.
Prompts of Subproblem Description Quality Validation is shown as follows:
Subproblem Description Quality Validation You are an expert evaluator of programming task subproblems. You will receive a subproblem (including the statement, solution code, and test cases). Evaluate whether the statement adequately meets each of the following five criteria. For each criterion, assign **1** if it is satisfied, or **0** if it is not. The total possible score is thus 5. 1. Clarity: Is the statement smoothly and understandably expressed, allowing a reader to grasp the task goal quickly? 2. Completeness: Does the description include all key elements needed to accomplish the subtask? 3. Accuracy: Is the description free of ambiguity or logical errors, and does it match the problem requirements and the provided code? 4. Feasibility: Can an engineer unambiguously determine and implement the required functionality—and pass functional tests—based solely on this description? 5. Professionalism: Does it use accurate, domain-appropriate terminology and a style fitting technical norms of coding tasks? Return your result **only** as a JSON dictionary with these five keys and values of 0 or 1. Please do not be overly strict. Assign a score of 1 to a criterion if it is even partially satisfied, allowing for minor imperfections. Only give a score of 0 if there are major or fundamental issues. Furthermore, ensure that the overall score for any subproblem is not lower than 2. Here is the subproblem:content Only return the JSON dictionary and remove the “‘json note, nothing else.
Table 4: Robustness of Subproblem Descriptions (Ref: Deepseek-V3). Left: Semantic Consistency; Right: Stylistic Similarity.
Model Semantic Stylistic
ROUGE-1 BERT Len. Ratio Marker Diff. Error Diff.
GPT-4o-mini 0.582 0.725 0.975 0.026 0.367
Gemini-2.0-Flash 0.600 0.734 1.141 0.019 0.728
o3-mini 0.554 0.692 1.181 0.124 1.350
Table 5: Structural Complexity Metrics Across Models
Model Avg. Turns Avg. Depth
o3-mini 2.08 1.94
Claude-3.7-Sonnet 2.08 1.94
Gemini-2.5-Pro 2.08 1.94
Table 6: Human Evaluation of Subproblem Quality. Metrics include Clarity (Cla.), Completeness (Com.), Accuracy (Acc.), Feasibility (Fea.), and Professionalism (Pro.).
Source Avg. Rating Cla. Com. Acc. Fea. Pro.
o3-mini 4.222 0.981 0.604 0.955 0.797 0.986
Claude-3.7 4.038 0.889 0.745 0.745 0.668 0.990
Human 4.202 0.981 0.659 0.904 0.726 0.933
A.7 Compute Cost Estimation
The cost of using LLM in our data curation pipeline is estimated as below:
  • Generating Solution Code: On average, each problem requires approximately 3K input tokens and 5K output tokens (including reasoning). The official Deepseek-R1 API is currently 4 CNY per 1M input tokens and 16 CNY per 1M output tokens. Therefore, the total cost for 5K+ problems is approximately 450 CNY.
  • Generating Subproblem Description: On average, each problem requires approximately 3K input tokens and 1K output tokens. The official Deepseek-V3 API is currently 2 CNY per 1M input tokens and 8 CNY per 1M output tokens. Therefore, the total cost for 5K+ problems is approximately 70 CNY.
The cost of processing the initial batch of 5K+ problems is manageable and also a one-time effort that does not need to be repeated. Future updates to the benchmark will involve only newly published problems and will therefore be much smaller in scale and less resource-intensive.
Appendix B CodeFlowBench-Comp Dataset
Fig. 12 demonstrates the statistical information of CodeFlowBench-main. Regarding the distribution, the data shows a clear preference for multi-step logic. The overall-turns histogram peaks at 2 turns (accounting for $\sim$2,100 problems), significantly outnumbering single-turn instances ($\sim$1,500), and exhibits a long-tail pattern extending to $\geq$7 turns. Similarly, the overall-depth distribution is heavily concentrated at Depth 2 (>>2,700 problems), indicating that nested dependency chains are the norm.
Furthermore, correlating these metrics with Codeforces difficulty ratings confirms their validity. As illustrated in the rating curves, we observe a strong positive correlation beyond the structural baseline: as the interaction turns and dependency depth increase, the average problem rating rises sharply—approaching 2,900 for tasks with $\geq$7 turns. This demonstrates that our structural metrics effectively capture the intrinsic complexity of logical problems.
(a) Distributions of overall-turns and overall-depth.
(b) The Correlations with Rating Levels.
Figure 12: Statistics of the overall-turns and overall-depth metrics in CodeFlowBench-Comp. Subfigure (b) shows inflection points at turns = 1 and depth = 1. This is attributed to the fact that competition-level problems are not restricted to multi-turn or deeply nested structures. Partial difficult problems are designed to be solvable by single function such as number theory related problems.
Appendix C CodeFlowBench-Repo Dataset
In this section, we provide a detailed overview of the CodeFlowBench-Repo dataset.
C.1 DomainEval
Overview of DomainEval
DomainEval is an auto-constructed benchmark specifically designed to evaluate LLMs’ coding capabilities across diverse real-world domains. Unlike general-purpose benchmarks that focus on common algorithmic tasks, DomainEval targets six distinct domains: Computation, Network, Basic, System, Visualization, and Cryptography. The dataset is curated from representative open-source GitHub repositories that possess high community recognition (typically over 100 stars) to ensure code quality.
Structurally, each subject in DomainEval consists of three essential components: a natural language instruction, a reference solution (target function), and a suite of valid test cases extracted via a test-method matching strategy. Crucially, to ensure executability, DomainEval preserves the necessary context required to run the target code. This design ensures that the tasks are not isolated snippets but are deeply embedded in the dependency structures of authentic software projects.
Rationale for Source Selection.
The selection of DomainEval as the foundation for CodeFlowBench-Domain is a deliberate design choice driven by two key factors:
1. Authenticity across Specialized Domains. First, DomainEval provides a rigorous source of ecologically valid code samples. By extracting problems from mature, production-level repositories, it captures the authentic coding patterns that professional developers actually employ. Furthermore, its explicit coverage of specialized domains like Cryptography and System areas where general LLMs often struggle allows our benchmark to assess Dependency Awareness in contexts that require precise API knowledge, effectively complementing the algorithmic logic focus of standard competition datasets.
2. Alignment with Constructive Codeflow Paradigm.
While many repository-level benchmarks focus on maintenance tasks such as issue resolution or debugging within existing codebases, CodeFlowBench aims to evaluate the constructive aspect of software engineering, namely to build complex functionality from the ground up. DomainEval focuses on function generation within a rich context. This format is the ideal input for our pipeline, allowing us to decompose a monolithic, dependency-heavy function into a step-by-step, iterative construction process. By leveraging DomainEval, we effectively demonstrate that the iterative codeflow paradigm is adaptable to building functional modules in authentic software libraries, validating our pipeline’s effectiveness beyond algorithmic puzzles.
C.2 Dataset Construction
The construction of the CodeFlowBench-Repo subset adapts the Subproblem Generation Phase from the CodeFlowBench-Comp pipeline. A example of problems in DomainEval is shown in Fig. 13
Figure 13: An example problem in DomainEval. We use the "method_code" attribute to generate subproblem solution code and use test case in "test_code" to generate subproblems’ test cases.
Since DomainEval already provides high-quality reference functions as ground truth, we omit the verification step and proceed directly to dependency analysis. We remove problems which only include single-turn function from the dataset and for the remaining, we parse the dependency relationship of the reference code to identify internal function calls and decompose them into iterative subproblems. Similarly, we then call LLM to generate problem description for subproblems. Finally, we use test cases of the top-level function to generate test cases of subproblems. An example is shown in Fig. 14.
Figure 14: An example problem in CodeFlowBench-Repo, which has a shared form of problems in CodeFlowBench-Repo.
To ensure robust evaluation, we parse the import statements from the original test files and utilize Docker containers to encapsulate the complex package dependencies required by these real-world repositories. This ensures a compatible and reproducible execution environment for all generated subproblems.
C.3 Dataset Statistics
Since our pipeline explicitly looks for internal function reuse to create multi-turn steps, the API-centric nature of the Computation domain resulted in empty dependency trees. Consequently, our selection was restricted to the remaining 749 problems from the other five domains. After further filtering, only 70 high-quality multi-turn problems survived the pipeline. To ensure a balanced evaluation, we augmented this set by add 12 single-turn problems each domain and curate the final CodeFlowBench-Repo.
As illustrated in Fig. 15, the dataset features a balanced distribution across five distinct domains: System (28.7%) and Cryptography (23.3%) constitute the majority, followed by Network (21.7%), Basic utilities (17.8%), and Visualization (8.5%). This distribution ensures the benchmark covers a wide spectrum of software engineering scenarios, from low-level OS interactions to high-level data plotting.
Structurally, the dataset presents significant complexity beyond atomic code generation. The distribution of overall turns and overall depths is shown in Fig. 15. The distribution of overall turns reveals that 56.6% of the problems require multi-turn iterative generation, forcing models to maintain reasoning continuity over long horizons. Furthermore, the dependency depth distribution confirms the intricate nature of these tasks: while 45.7% of problems act as foundational nodes (Depth 1), the remaining 54.3% involve deeper dependency chains (Depth 2–4), requiring the model to correctly implement and call prerequisite functions within a hierarchical context.
Figure 15: Statistics of CodeFlowBench-Repo. Left: The diversity of domains, with a focus on System and Cryptography tasks. Right: Distributions of interaction turns and dependency depths, highlighting the structural complexity.
Appendix D Mathematical Expression of PD@kPD@k& $APD@k$
For a given problem, we define $PD@k$ as the expected maximum pass depth over $k$ independent trials of the model. Directly using only those $k$ results leads to high variance, just as with the $pass@k$ metric. By analogy to the unbiased estimator for $pass@k$, which leverages $n$ trials ($n>k$) to reduce variance, we derive a similar estimator for $PD@k$.
Let the pass depths from $n$ trials be:
$\{d_{1},d_{2},\dots,d_{n}\},$
and let
$d_{(1)}\leq d_{(2)}\leq\cdots\leq d_{(n)}$
denote these depths in ascending order. Then an unbiased estimator for $PD@k$ is
$PD@k\;=\;\sum_{j=k}^{n}d_{(j)}\,\frac{\binom{j-1}{k-1}}{\binom{n}{k}}$. (1)
To see how this arises, consider sampling a random subset of size $k$ from the $n$ depths. Let $M\;=\;\max\{\,d_{i_{1}},d_{i_{2}},\dots,d_{i_{k}}\}$ be the maximum depth in that subset. An unbiased estimator of $M$ is $\mathbb{E}[M]$. By construct $\mathbb{E}[M]\;=\;PD@k$ ,this estimator is unbiased since its expected value exactly equals the true expected maximum depth. Consider all possible value of $M$, we have:
$\mathbb{E}[M]=\sum_{m}m\,P(M=m)=\sum_{j=k}^{n}d_{(j)}\,P\bigl(M=d_{(j)}\bigr).$ (2)
The probability that $M=d_{(j)}$ equals the probability of choosing $d_{(j)}$ along with k1k-1 depths from the first j1j-1 smaller values:
$P\bigl(M=d_{(j)}\bigr)=\frac{\displaystyle\binom{j-1}{k-1}}{\displaystyle\binom{n}{k}}$. (3)
Combining (3) and (4) immediately recovers (1).
Finally, we define
$APD@k=\frac{1}{|\mathcal{P}|}\sum_{p\in\mathcal{P}}PD@k(p)$ (4)
the average $PD@k$ over the set $\mathcal{P}$ of all problems.
Appendix E Experiment
E.1 Experiment Setting
Due to the substantial size of the total question pool ($N=5258$), we implemented stratified sampling with proportional allocation across overall-depth categories to select 1,000 test samples, as detailed in Table 7. To validate sampling quality, we conducted a $\chi^{2}$ test comparing the overall-turn distributions between the population and sampled data (Table 8). The statistical analysis yielded a p-value of 0.1246, indicating no significant difference ($\alpha=0.05$) in distribution patterns. This confirms the representativeness of our sampling strategy and ensures the validity of subsequent analytical outcomes.
Table 7: Comparison of Overall-Depth Proportions Between the Population and the Sample
Overall-Depth # Population Pop. Proportion (%) # Sample Sam. Proportion (%)
1 1,488 28.3 283 28.3
2 2,751 52.3 523 52.3
3 870 16.5 165 16.5
4 125 2.4 24 2.4
$\geq$5 24 0.5 5 0.5
Total 5,258 100.0 1,000 100.0
Table 8: Comparison of Overall-Turn Proportions Between the Population and the Sample
Overall-Turn # Population Pop. Proportion (%) # Sample Sam. Proportion (%)
1 1,488 28.3 283 28.3
2 2,158 41.0 437 43.7
3 990 18.8 188 18.8
4 402 7.6 58 5.8
5 149 2.8 21 2.1
6 46 0.9 6 0.6
$\geq$7 25 0.5 7 0.7
Total 5,258 100.0 1,000 100.0
E.2 Instruction Templates
In the instruction template for multi-turn testing, we implemented a set of basic heuristics to adapt to different problem types. For example, if a question has no dependencies, we omit the “## Dependency information” section. If no prior code is provided—which is common when the overall depth is 1—we exclude the “## Pre-implemented functions” section. Furthermore, if it is the final turn of the code, we append the following snippet:
import sys
def {name}():
input = sys.stdin.read().split()
Similarly, if the problem does not have any dependencies, we will also omit the section related to {dependencies}.
Multi-turn Test You are a Programming Expert. You always provide correct and reliable code solutions. You will be provided with the background of the whole problem, a programming problem and may also some pre-implemented functions.If pre-implemented functions provided, you need to call the pre-implemented functions and write a new function to solve the problem. ## Background of the whole problem: {problem_description} ## Problem Description: You need to complete name function. {statement} ## Dependency information: To solve the problem, you need to utilize the ## Pre-implemented functions {dependencies} provided. ## Pre-implemented functions: {history} ## Guidelines: - Ensure the function is executable and meets the requirement. - Handle ## Dependency information correctly. - Provide clear and concise comments to explain key parts of the code. Return your response by filling the function body following the function signature provided. Just generate the function and don’t output any examples.
[…]
Appendix F Discussion
F.1 Discussion on Models’ DSC Performance
For a problem, the DSC metric is defined as:
$\text{DSC(problem)}=\frac{\text{Overall-Turns(problem)}}{\text{Overall-Depth(problem)}}$
Recall that the overall-turn and overall-depth of a problem are derived from its AST, corresponding to the number of nodes and the depth of its AST. Based on this, we can see that a high DSC value indicates a problem with a complex dependency structure. Figure 16 presents the pass@1 scores of models across different DSC intervals. It can be observed that most models are only capable of solving problems with DSC equal to 1, which corresponds to a simple linear dependency structure. Only a few leading models are able to solve a limited number of problems with DSC values below 1.33. All models struggle significantly when faced with problems involving more complex structures.
Figure 16: Heatmap of models’ pass@1 scores on multi-turn problems within different DSC intervals.
F.2 Error Case Study
We have defined three typical error types for test in multi-turn pattern, in this part we will introduce several example to illustrate them.We have simplify and arrange model’s outputs to make it clear to read. The content of "generated" field is model’s output. The content of "harness_result" field is the verification result by running it with test cases."1" denotes accepted, "0" denotes wrong answer and "wrong" denotes running error.
Figure 17: Example of an Incomplete Reasoning (IR) Error by Deepseek-V3.The original problem is 1799F.The algorithm produced by the model fails to find the global optimum because it greedily sorts each element according to the single-step gain—choosing either the halving operation or the subtraction operation in isolation. In reality, achieving the global optimum requires a two-dimensional dynamic-programming solution that considers both operations jointly.
Figure 18: Example 1 of an Insufficient Globalization(IG) error by Deepseek-V3.The original problem is 2005D.The model’s generated code omitted the import math statement, resulting in an error when calling gcd. This issue stems from improper handling of external imports.
Figure 19: Example 2 of an Insufficient Globalization(IG) error by Deepseek-V3.The original problem is 1734F Because the program reads new inputs and performs fresh calculations for each test case, failing to clear the cache beforehand can inadvertently reuse stale results from a previous case. This leads to incorrect output due to unintended interaction between caching and I/O.
Figure 20: Example of an Instruction Misinterpretation(IM) error by Deepseek-V3.The original problem is 1575H Although the DP routine correctly unpacks both transition and match_counts from build_transition(b), it merely treats match_counts as an ordinary list. In reality, match_counts carries two essential pieces of semantic information.(1)Match indicator.It denotes a complete match of b.(2)Backtrack hint. It’s a hint in combination with the prefix function, indicating how far the automaton should jump back after a match to continue detecting overlapping occurrences. This error is caused by models’ misinterpretation to the dependency relationship between tool function and top-level function
Appendix G Impact Statement
This paper introduces CodeFlowBench, a comprehensive benchmark for evaluating code generation models in realistic multi-turn, dependency-driven development scenarios. For research, CodeFlowBench fills a critical gap by providing a standardized suite of tasks that require iterative reasoning, function dependency management, and end-to-end solution assembly. By exposing models’ deficiencies in global awareness, instruction consistency, and dependency handling, CodeFlowBench will catalyze the design of new architectures and training paradigms that explicitly model iterative workflows and cross-turn coherence. Its open dataset and evaluation protocol invite the community to develop and compare dependency-aware generation strategies, driving progress toward more robust and developer-friendly code assistants.
In industry, CodeFlowBench offers a practical yardstick for assessing the readiness of AI coding tools in real-world software development, where tasks rarely appear as isolated single-step prompts. Integrating CodeFlowBench into CI/CD pipelines can help organizations detect and remediate weaknesses in model-powered code suggestions before deployment, reducing debugging overhead and technical debt. By highlighting the importance of function reuse, import management, and state consistency across revisions, CodeFlowBench insights can inform best practices for AI-augmented coding workflows, accelerating adoption of reliable co-programming solutions. There are broader societal implications in enabling safer, more maintainable AI-generated code, yet none that we believe warrant special emphasis here.